# Stochastic Block Covariance Matrix Estimation

### v1

@@ 6.2 Monthly stock returns of energy and financial companies
Estimating the covariance matrix is a critical aspect of the financial market, particularly in the realms of portfolio optimization and risk management. Covariance, which gauges the co-movements of returns across diverse assets, forms the basis for comprehending the relationships between these assets. Through an accurate covariance-based classification, investors can gain insights into the interconnections among different assets, facilitating portfolio diversification to mitigate specific risks.
To underscore the importance of covariance structure, we reference a concise example from Liechty et al. (2004). In this example, we examine monthly stock returns spanning April 1996 to December 2001 for nine companies (68 time points). These companies belong to either the energy sector (Reliant, Chevron, British Petroleum, and Exxon) or the finance sector (Citi-Bank, Lehman Brothers, Merrill Lynch, and Bank of America). Notably, Enron underwent a transformation from an energy company to a finance company in the 1990s and subsequently declared bankruptcy in December 2001. In our analysis, we focus solely on data prior to Entron’s bankruptcy, differing from the analysis presented in Liechty et al. (2004).
Figure 9: Monthly returns of Enron and other energy and finance companies.
We aim to: (1) cluster these nine companies to assess whether their equity variations align with their industrial classifications, and (2) evaluate Enron’s transition from the energy sector to the finance sector prior to its bankruptcy. Given the significant fluctuations in some stocks, assuming a block structure in the correlation matrix is more appropriate. Therefore, we apply our estimator to the standardized data .
Figure 9 displays a trace plot of standardized monthly returns, presenting a challenge to discern whether Enron’s fluctuations are more in sync with finance or energy companies. The block covariance estimation, shown in Figure 10, reveals three main clusters: Enron and Reliant (energy), other energy companies, and finance companies. These classification align with their respective industrial classifications.
Figure 10: Estimators of covariance matrix (left: sample covariance, right: block covariance under the hierarchical prior) and Posterior similarity matrix. Enron behaves more closely resembles that of energy companies than finance companies.
It appears that Enron did not successfully transition from the energy sector to finance, as its behavior more closely resembles that of energy companies than finance companies. Enron and Reliant share similarities, exhibiting lower covariance with other energy companies and finance companies compared to the remaining energy companies. This finding is consistent with the conclusion of Liechty et al. (2004), which also indicated Entron’ unsuccessful transition.
@@ 6.3 Phenotypic integration of Mediterranean shrubs under contrasting environmental conditions
The analysis of correlation structures among traits provides insight into phynotypic integration, which refers to "the pattern of functional, developmental, and/or genetic correlation among different traits in a given organism" (Pigliucci, 2003). Traits may be positively or negatively correlated depending on their functional roles, influencing how organisms adapt to varying environmental conditions. To examine phenotypic integration under different environmental conditions, we analyze the correlation matrices of 20 ecophysiological traits in Mediterranean shrub samples (Lepidium subulatum L., Brassicaceae) grown under both stressful (drought) and favorable conditions. This dataset allows us to assess differences in phenotypic integration between these environments. The experimental details are described in Matesanz et al. (2021).
The dataset is aggregated at the family level, where we computed the average trait values for each population and maternal family. After removing missing cases, we obtained 45 samples from plants grown under drought conditions and 55 from those in favorable conditions. To better meet the normality assumptions, we applied logarithmic or square-root transformations to most traits before standardization. The 20 traits can be categorized into the following functional groups:
@@
These traits may either compete for resources or work together in a coordinated manner.
@@
Figure 11: Posterior mean of covariance matrices of traits under favorable and stressful environments. The overall block structure aligns well with functional group classifications, with notable differences in correlation patterns across environments. Under drought conditions, cross-functional correlations strengthen, suggesting an adaptive response to environmental stress.
@@
Figure 11 presents the sample correlation matrices and their posterior mean estimates under favorable and stressful conditions. The overall block structure aligns well with functional group classifications. Notably, reproductive traits (IS, FN, IN, RO) exhibit strong positive correlations, reflecting coordinated investment in reproduction. Similarly, phenological traits (FR, FBF, FL) form a highly correlated cluster. These traits occur in sequence during the reproductive timeline, which explains their strong correlation. The morphological traits (LA, LL, PH, TELA) are positively correlated with above-ground biomass (AB), indicating an integrated response to plant growth. Additionally, a cross-functional cluster emerges, incorporating physiological (RGR, FvFm), allocation (R.L, TB), phenological (Sen), and reproductive (SS) traits.
@@
Under drought conditions, the correlation matrix exhibits enhanced correlation across functional groups compared to the favorable environment. (1) Stronger negative correlations between reproductive phenology (FBF, FL, FR) and other traits. (2) Greater positive correlations between reproductive traits (IS, FN, IN, RO) and traits in the cross-functional group (Sen, R.L, TB, FvFm, SS). (3) Enhanced positive correlations between above-ground growth traits (LA, LL, PH, TELA, AB) and traits in the cross-functional group (Sen, R.L, TB, FvFm, SS).
@@
This pattern is consistent with a drought-escape strategy (Blanco-Sánchez et al., 2022; Matesanz et al., 2020), wherein plants accelerate reproductive timing and enhance growth to complete life cycle before drought conditions become more severe. For example, a larger cross-functional cluster emerge, involving coordinated changes in physiological, phenological, and allocation traits. Under drought conditions, plants that begin reproduction earlier tend to grow larger (higher PH, TB, LA), produce larger seeds, show increased leaf senescence, and exhibit higher growth rates and photochemical efficiency.
@@
Matesanz et al. (2021) quantified phenotypic integration in plants by counting the number of significantly nonzero pairwise correlations between traits and found that plants under stressful conditions exhibit higher phenotypic integration. In our model, this increased integration appears as a reduction in the number of distinct trait clusters (see posterior similarity matrix in Figure 19 for a clear block assignment). Under drought conditions, traits from different functional groups that are only weakly or negatively associated in favorable conditions become more strongly or positively associated, suggesting an adaptive reconfiguration of trait relationships to cope with environmental stress.
@@ 7 Discussion
We introduced a stochastic block structure for covariance matrix estimation, motivated by practical demands in real-world applications. Beyond its interpretability, the stochastic block structure effectively reduces dimensionality, while being more general, flexible, and efficient compared to conventional sparsity assumptions or shrinkage estimators. The block structure also offers greater adaptability across different types of covariance matrices, enhancing its utility in various fields.
@@
In many practical applications, the block structure may be more relevant for the correlation matrix than the covariance matrix. For example, principle component analysis and copula models rely on correlation matrices. In portfolio selection, some assets may exhibit large variances, which can dominate clustering results and obscure important co-movements. When clustering is performed based on the covariance matrix, variables with large variances often form distinct blocks, pushing variables with smaller variances into a single block. This sensitivity to variance can lead to suboptimal clustering that fails to capture meaningful patterns of co-fluctuation.
@@
To more accurately capture the block structure in a correlation matrix, we recommend standardizing the data to have zero mean and unit variance before applying the block covariance model. However, standardizing data is not always feasible, as block correlation matrices may be an intermediate product rather than the target estimator. Furthermore, standardization can ignore the uncertainty quantification of variance estimation by treating variance as fixed. A more direct approach is to consider a canonical representation of block correlation matrices. However, this introduces computational challenges due to the inherent restriction that correlation $r_{ij}\in[-1,1]$, which imposes constraints on the parameters $(A,\mathbf{\lambda})$. These constraints make it difficult to apply standard techniques, as a conjugate prior is not readily available, complicating both the estimation of correlations and the block assignment process.
A better solution might involve introducing scaling parameters to account for heteroskedasticity. We can express the covariance matrix as $\Sigma=S^{1/2}\Sigma S^{1/2}$, where $S$ is a diagonal matrix representing individual variances, and $\Sigma$ is a block covariance matrix.
@@
One of the main challenges of our method is the computational burden, especially for large datasets where the number of variables $p$ exceeds 200, as often encountered in proteomics, gene expression, and finance. The bottleneck lies in estimating the block assignments, where even with an effective merge-split sampler (Dahl and Newcomb, 2022), we encounter slow mixing and high computational costs.
Several strategies could improve computational efficiency. One solution is to enhance the Metropolis Hasting proposal in the sampler. Currently, the sampler is constrained to move between $K+1$ (split) and $K-1$ (merge) clusters. For $K-1$ cases, there is only one way to merge, while for $K+1$ cases, the splitting depends on the previous allocated items, which may be inefficient. A more flexible proposal could be constructed by leveraging clustering structures from previous samples, potentially based on the posterior similarity matrix. Another approach, inspired by Duan and Bhattacharya (2024), is to construct a graph where each node represents a potential state, and then perform MCMC using random walks over the graph. By optimizing the graph’s nodes and edges, larger jumps toward more likely structures can be achieved, improving the efficiency of the sampling process. However, designing this graph and understanding its theoretical properties require careful consideration, as the effectiveness of this method depends on how well the graph structure reflects the underlying block structure and ensures adequate exploration of the posterior distribution.
Instead of focusing on improving the sampler itself, we could explore parallelizing the sampling process to alleviate the computational burden. A well-known approach involves divide-and-conquer (D&C) strategies (Sivaganesan et al., 2017; Chang and Fisher, 2013; Ni et al., 2019; Lovell et al., 2018; Srivastava et al., 2015; Campbell and Broderick, 2019), which divide the problem into simpler, independent components that are processed in parallel, and then combine the results to approximate the full solution. Unlike the traditional D&C methods, which break down large-scale data problems into small datasets, we break down the variables themselves, with some variables appearing across multiple subsets to serve as benchmarks. Each subset of data can be modeled independently, and the posterior samples can be "sutured" together based on the clustering information provided by these benchmark variables. Determining how to appropriately divide the variables and merge the posterior samples to obtain a reliable and effective approximation of the target estimator requires careful planning and methodological considerations.
@@ F.2 Extra figures from applications
Figure 16: Covariance estimation of neural population coordination under different priors (sorted by block assignments of hierarchical prior)
Figure 17: Posterior distribution of number of blocks under different priors in the mutual funds application.
Figure 18: Posterior similarity matrix under different priors in the mutual funds application.

### v2

@@ 6.2 Monthly stock returns of energy and financial companies
Estimating the covariance matrix is crucial in the financial market, particularly for portfolio optimization and risk management. An accurate covariance-based classification can provide insights into the interconnections among different assets, facilitating portfolio diversification to mitigate specific risks.
To illustrate covariance-based classification, we revisited the example from Liechty et al. (2004) but addressed in a flexible way without prefixing the number of groups. In this example, we examined monthly stock returns spanning April 1996 to December 2001 for 9 companies (68 time points). These companies belonged to either energy sector (Reliant, Chevron, British Petroleum, and Exxon) or finance sector (Citi-Bank, Lehman Brothers, Merrill Lynch, and Bank of America). We only focused on data prior to Enron’s bankruptcy (December 2001), differing from the analysis presented in Liechty et al. (2004).
Figure 7: Monthly returns of Enron and other energy and finance companies. It was unclear whether Enron’s fluctuations were more in sync with finance or energy companies
We aimed to: (1) cluster these 9 companies to assess whether their variations align with their industrial classifications, and (2) evaluate Enron’s transition from the energy sector to the finance sector prior to its bankruptcy. Given the significant fluctuations in some stocks, we assumed group covariance structure in the correlation matrix, applying our estimator to the standardized data.
Figure 7 displays a trace plot of standardized monthly returns, presenting a challenge to discern whether Enron’s fluctuations were more in sync with finance or energy companies. The stochastic block covariance estimation, shown in Figure 8, revealed three main clusters: Enron and Reliant (energy), other energy companies, and finance companies. These classification aligned with their respective industrial classifications.
Figure 8: Estimators of covariance matrix (left: sample covariance, right: stochastic block covariance estimation under the hierarchical prior) and posterior similarity matrix. Enron behaved more like energy companies than finance companies.
It appeared that Enron did not successfully transit from the energy to finance sector, as it behaved more closely resembled energy companies than finance companies. Enron and Reliant shared similarities, exhibiting lower covariance with other energy companies and finance companies compared to the remaining energy companies. This finding was consistent with the conclusion of Liechty et al. (2004), which also indicated Entron’ unsuccessful transition.
@@ 6.3 Phenotypic integration of Mediterranean shrubs under contrasting environmental conditions
Analyses of correlation structures among traits can offer insights into phenotypic integration, i.e., “the pattern of functional, developmental, and/or genetic correlation among different traits in a given organism” (Pigliucci, 2003). Traits may be positively or negatively correlated depending on their functional roles, influencing how organisms adapt to varying environmental conditions. To examine phenotypic integration under different environmental conditions, we present here an analysis of the correlation matrices of 20 ecophysiological traits in Mediterranean shrub samples (Lepidium subulatum L., Brassicaceae) grown under both stressful (drought) and favorable conditions with a focus on assessing the differences in phenotypic integration between these environments. The experimental details are described in Matesanz et al. (2021).
The dataset recorded traits of 355 individual plants across 4 populations, with around 15 to 20 families nested within each population. We aggregated plant-level data to the family-level by computing the average trait values for each maternal family within each population. After removing missing cases, we obtained 45 samples from plants grown under drought conditions and 55 from those in favorable conditions. To better meet the normality assumptions, we applied a square-root transformation to two traits (FN, PH), and a logarithmic transformation to eight traits (FBF, FL, IN, LA, LL, RO, TB, TELA) before standardization. These 20 traits could be categorized into the following functional groups:
@@
These traits might either compete for resources or work together in a coordinated manner.
@@
Figure 9: Posterior mean of covariance matrices of traits under favorable and stressful environments. The overall block structure aligned well with functional group classifications, with notable differences in correlation patterns across environments (A1, A2, A3, A4). Under drought conditions, cross-functional correlations strengthened, suggesting an adaptive response to environmental stress (B, C, D).
@@
Figure 9 presents the sample correlation matrices and their posterior mean estimates under favorable and stressful conditions. The overall block structure aliged well with functional group classifications. Notably, reproductive traits (IS, FN, IN, RO) exhibited strong positive correlations, reflecting coordinated investment in reproduction (Figure 9 A1). Similarly, phenological traits (FR, FBF, FL) formed a highly correlated cluster (Figure 9 A2). These traits occurred in sequence during the reproductive timeline, which explained their strong correlation. The morphological traits (LA, LL, PH, TELA) were positively correlated with above-ground biomass (AB), indicating an integrated response to plant growth (Figure 9 A3). Additionally, a cross-functional cluster emerged, incorporating physiological (RGR, FvFm), allocation (R.L, TB), phenological (Sen), and reproductive (SS) traits (Figure 9 A4).
@@
Under drought conditions, the correlation matrix exhibited enhanced correlation across functional groups compared to the favorable environment: (1) stronger negative correlations between reproductive phenology (FBF, FL, FR) and other traits (Figure 9 B); (2) greater positive correlations between reproductive traits (IS, FN, IN, RO) and traits in the cross-functional group (Sen, R.L, TB, FvFm, SS) (Figure 9 C); (3) enhanced positive correlations between above-ground growth traits (LA, LL, PH, TELA, AB) and traits in the cross-functional group (Sen, R.L, TB, FvFm, SS) (Figure 9 D).
@@
This pattern was consistent with a drought-escape strategy (Blanco-Sánchez et al., 2022; Matesanz et al., 2020), wherein plants accelerate reproductive timing and enhance growth to complete life cycle before drought conditions become more severe. For example, a larger cross-functional cluster emerged, involving coordinated changes in physiological, phenological, and allocation traits. Under drought conditions, plants that began reproduction earlier tended to grow larger (higher PH, TB, LA), produced larger seeds, showed increased leaf senescence, and exhibited higher growth rates and photochemical efficiency.
@@
Matesanz et al. (2021) quantified phenotypic integration in plants by counting the number of significantly non-zero pairwise correlations between traits and found that plants under stressful conditions exhibited higher phenotypic integration. In our model, this increased integration appeared as a reduction in the number of distinct trait clusters (see posterior similarity matrix in Figure 15 for a clear block assignment). Under drought conditions, traits from different functional groups that were only weakly or negatively associated in favorable conditions became more strongly or positively associated, suggesting an adaptive reconfiguration of trait relationships to cope with environmental stress.
@@ 7 Discussion
We introduced a stochastic block structure for covariance matrix estimation, motivated by practical demands in real-world applications. Beyond its interpretability, the stochastic block structure effectively reduces dimensionality, while being more general, flexible, and efficient compared to conventional sparsity assumptions or shrinkage estimators. The stochastic block structure also offers greater adaptability across covariance matrices with different structures, enhancing its utility in various fields.
@@
In many practical applications, the stochastic block structure may be more relevant for correlation matrices than covariance matrices. For example, principle component analysis and copula models rely on correlation matrices. In portfolio selection, some assets may exhibit large variances, which can dominate clustering results and obscure important co-movements. When clustering is performed based on the covariance matrix, variables with large variances often form distinct blocks, pushing variables with smaller variances into a single block. This sensitivity to variance can lead to suboptimal clustering that fails to capture meaningful patterns of co-fluctuation.
@@
To more accurately capture the block structure in a correlation matrix, we recommend standardizing the data to have zero mean and unit variance before applying the stochastic block covariance model. However, standardizing data is not always feasible, as correlation matrices may be an intermediate product rather than the target estimator. Furthermore, standardization can ignore the uncertainty quantification of variance estimation by treating variance as fixed. A more direct approach is to consider a semi-spectral decomposition of block correlation matrices. However, this introduces computational challenges due to the inherent restriction that correlations $r_{ij}\in[-1,1]$, imposing constraints on the parameters $(A,\mathbf{\lambda})$. These constraints make it difficult to apply standard techniques, as a conjugate prior is not readily available, complicating both the estimation of correlations and the block assignment process. A better solution might involve introducing scaling parameters to account for heteroskedasticity. We can express the covariance matrix as $\Gamma=V^{1/2}RV^{1/2}$, where $V$ is a diagonal matrix representing individual variances, and $\Sigma$ is a block covariance matrix.
@@
A substantial challenge to our method is the computational cost which scales poorly with large datasets. Our current implementation cannot handle datasets with $p>200$, as often encountered in proteomics, gene expression, and finance. The bottleneck lies in estimating the block assignments under partition priors, where even with an effective merge-split sampler (Dahl and Newcomb, 2022), we encounter slow mixing and high computational costs. It would be worthwhile to innovate more efficient computing strategies which can overcome the incremental nature of present Markov chain samplers in exploring the space of partitions.
8 Acknowledgement
We thank Marlene Cohen, Douglas Ruff, Silvia Matesanz, Micaela Carvajal, Jonathan Azose, and Jason Xu for sharing their data with us. We also thank Jaeho Kim and Drew Creal for providing their code and detailed numerical experiment settings. We benefitted from helpful discussion with Ed Tam and Johan Vdmolen. The work was partially supported by NIH awards R01 DC013096 and R01 DC016363.